# xOffense: An AI-driven autonomous penetration testing framework with offensive knowledge-enhanced LLMs and multi agent systems

### v1

@@ 5.1.1 Environment
    AUTO-PEN-BENCH [10]: Tasks were instantiated as Docker containers, which were hosted on a separate virtual machine to avoid resource contention with the attacker host. This victim VM was configured with 4 vCPUs, 8 GB RAM, and 80 GB storage, and placed in the same NAT network as the attacker machine to ensure direct connectivity.
@@
    AI-Pentest-Benchmark [11]: Vulnerable machines were directly imported from official VulnHub distributions and executed as VMware virtual machines without modification to their default specifications, in order to preserve the original exploitation conditions. All machines were assigned to the same NAT network as the attacker host to guarantee consistent communication.
5.1.2 Model Serving
@@
The fine-tuned model, namely Qwen3-32B-finetune was hosted on a dedicated compute node equipped with an NVIDIA A100 GPU (80 GB VRAM). The same hardware was also used during the fine-tuning process to accelerate training efficiency. For inference, the model was exposed via an API endpoint tunneled through ngrok, allowing the attacker machine to interact with the model as an external service. This separation ensured that LLM inference did not compete with penetration testing tasks for system resources, while also replicating realistic deployment conditions where models are often served remotely.
5.1.3 Evaluated Models
The evaluation considered both commercial and open-source large language models to provide a comprehensive baseline. Specifically, we included GPT-4o, Llama3.3-70B, Llama3.1-405B, and DeepSeek-V3. For open-source mid-scale models, we assessed Qwen3-32B-base, which provides a strong balance between efficiency and capability. Building upon this foundation, we fine-tuned the model into Qwen3-32B-finetune, specialized for penetration testing tasks within xOffense. All models were evaluated with a temperature setting of 0.5 to balance deterministic output with generative flexibility, ensuring consistency in automated penetration testing scenarios.
5.2 Fine-Tuning Methodology
@@
Unlike conventional large-scale model adaptations, which demand significant computational overhead, xOffense leverages LoRA (Low-Rank Adaptation) to achieve domain specialization. LoRA reduces the parameter footprint by freezing the base model’s weights and training only a compact set of adapter matrices. This approach dramatically lowers the number of trainable parameters by over 99%, enabling efficient fine-tuning even on standard GPU infrastructures.
@@
To further address the memory bottlenecks of handling a 32B-parameter model, we employ DeepSpeed ZeRO-3 [42] optimization. ZeRO-3 partitions model states—including optimizer, gradients, and parameters—across multiple GPUs, achieving linear scalability. Additionally, FlashAttention v2 [43] is integrated to optimize attention computation, reducing memory usage and accelerating training by up to 3x compared to standard attention implementations. These combined techniques allow us to efficiently fine-tune Qwen3-32B for penetration testing workloads with a significant reduction in hardware demands. And the fine-tuning dataset comprised two main corpora as follows:
@@
    PentestData: This dataset is meticulously curated to encompass domain-specific question-answer pairs, each enriched with synthetically generated Chain-of-Thought (CoT) reasoning traces. The reasoning steps are encapsulated within <think> tags, enabling the model to learn structured, step-by-step logical deduction processes tailored for penetration testing scenarios. To construct PentestData, we aggregate and standardize write-ups from over 1,000 machines across leading cybersecurity platforms, including TryHackMe [44], HackTheBox [45], and VulnHub [46]. These write-ups are systematically processed to extract task-specific reasoning paths, exploit procedures, and decision-making sequences relevant to offensive security operations. In addition, we incorporate supplementary pentest datasets from HuggingFace Datasets Hub [47], focusing on cybersecurity knowledge bases, penetration testing techniques, and practical guides for commonly used security tools. This comprehensive integration ensures that PentestData serves as a robust and diverse resource for training models in autonomous penetration testing workflows.
@@
    WhiteRabbitNeo: A high-quality JSONL-formatted dataset comprising instruction-response pairs, specifically curated for cybersecurity tasks. Although the original dataset lacked explicit Chain-of-Thought (CoT) reasoning annotations, it was standardized during preprocessing by appending empty <think> tags to each sample. This structural unification ensures compatibility with CoT-augmented training pipelines and facilitates subsequent fine-tuning for step-by-step reasoning abilities. The dataset draws from real-world offensive and defensive cybersecurity scenarios, encompassing exploitation techniques, payload crafting, and red-team/blue-team interactions, sourced from the WhiteRabbitNeo community contributions [48].
@@ 5.3 Benchmark Datasets
To ensure a rigorous and practically relevant evaluation, we selected two complementary benchmarks that cover both synthetic vulnerabilities and realistic multi-stage exploitation scenarios. Together, these benchmarks provide a balanced testbed for assessing the adaptability and robustness of automated penetration testing systems.
@@
This benchmark defines a total of 33 penetration testing tasks, spanning both instructional “in-vitro” exercises and real-world CVEs. The in-vitro set (22 tasks) reflects fundamental vulnerability classes frequently highlighted in industry rankings such as the OWASP Top 10, including weak access control (e.g., misconfigured sudo, world-writable shadow files), web application flaws (e.g., path traversal, SQL injection, file upload RCE), and insecure network configurations (e.g., SNMP misconfiguration, ARP spoofing). In addition, four cryptography tasks evaluate resilience against improper or weak cryptographic implementations, such as brute-forcing Diffie–Hellman keys. Beyond these educational tasks, the benchmark incorporates 11 real-world CVEs ranging from 2014 to 2024 with CVSS scores between 7.5 and 10.0. These include critical vulnerabilities widely recognized for their impact and prevalence, such as Log4Shell (CVE-2021-44228), Heartbleed (CVE-2014-0160), SambaCry (CVE-2017-7494), and Spring4Shell (CVE-2022-22965). By combining foundational categories with critical CVEs, AUTOPENBENCH provides a structured yet realistic environment to evaluate whether agents can transition from basic exploitation to handling high-severity vulnerabilities that have historically dominated real-world attack campaigns.
@@
While AutoPenBench focuses on containerized vulnerabilities, the AI-Pentest-Benchmark evaluates AI agents on complete end-to-end exploitation workflows across 13 real-world vulnerable machines drawn from VulnHub. These machines are categorized by difficulty into easy (e.g., Victim1, Library2, Funbox, WestWild), medium (e.g., Cengbox2, Devguru, Symfonos2), and hard (e.g., Insanity, TempusFugit). Each machine defines a structured set of reconnaissance, exploitation, privilege escalation, and general technique subtasks, amounting to 152 tasks in total. The vulnerabilities embedded within these machines reflect common penetration testing scenarios, including web application flaws (SQL injection, XSS, CSRF/SSRF), network service weaknesses (FTP/AD enumeration, brute-force authentication), code-level issues (deserialization, command injection), and post-exploitation techniques (cronjob analysis, misconfigured system files, privilege escalation via user access exploitation). Notably, many of these tasks map directly to recurring weakness categories in the CWE Top 25 and OWASP Top 10, ensuring that success on this benchmark corresponds to capabilities relevant in practical offensive security operations. The benchmark is particularly challenging because the ultimate goal is to achieve root access on each machine, requiring coherent reasoning across reconnaissance, exploitation, and privilege escalation stages. Previous studies have shown that even large-scale proprietary models such as GPT-4o and Llama3.1-405B were unable to achieve root-level compromise without human assistance, underscoring the difficulty and realism of this benchmark.
@@ 5.4.1 Overall Task Completion Rate
This metric measures the percentage of target machines successfully compromised (i.e., the flag was obtained) within the allowed interaction budget. Formally:
@@ 5.4.2 Sub-task Completion Rate (1 Experiment)
To gain insight into intermediate stages of penetration testing, we evaluate sub-task success rates (e.g., service enumeration, vulnerability detection). Each benchmark defines a set of subtasks $S$. A sub-task is considered successful if it is completed in at least one of the five independent runs:
@@ 6.1 Experiment Scenarios
6.2 Task Completion Performance Across Penetration Testing Categories
@@ 6.2.1 Overall Task Completion Performance
Table 2 presents the overall task completion rates across all evaluated models on the AUTO-PEN-BENCH dataset. The fine-tuned Qwen3-32B-finetune model achieved a remarkable 72.72% completion rate, substantially outperforming both its base variant, Qwen3-32B-base (30.30%), and other state-of-the-art models, including GPT-4o (21.21%), Llama3.1-405B (Paper) (30.30%), and PentestGPT (9.09%).
The performance disparity between Qwen3-32B-finetune and its base version is particularly noteworthy. Despite having identical model architecture and parameter size (32 billion parameters), the domain-specific fine-tuning enabled a 2.4x improvement in task completion. This validates the effectiveness of our lightweight LoRA fine-tuning pipeline in adapting general-purpose models to specialized penetration testing workflows.
In the Access Control (AC) category, Qwen3-32B-finetune achieved a 100% success rate, a stark contrast to the 40.00% of Qwen3-32B-base and the 60.00% of Llama3.1-405B (Paper). Similarly, in Network Security (NS), our model achieved 83.33% completion, surpassing all baselines, including Llama3.3-70B (33.33%) and Qwen3-32B-base (50.00%). Notably, even in the complex Real-world category, Qwen3-32B-finetune attained a 54.54% success rate, outperforming Qwen3-32B-base (27.27%) and PentestGPT (0.00%).
These findings demonstrate that fine-tuning on domain-relevant CoT data and incorporating robust task orchestration mechanisms can enable a quantized, resource-efficient model to match—and even surpass—the capabilities of larger, general-purpose LLMs in specialized scenarios. The consistent outperformance across categories further validates the robustness of our fine-tuning strategy, especially given the compute-efficient AWQ quantization.
@@ 6.2.2 Sub-task Completion Performance (1 Experiment)
To assess finer-grained capabilities, we evaluated sub-task completion in a single-run experiment (Table 3). The term ”1 Experiment” refers to the overall subtask completion rate across five experiments, where a subtask is considered successful if it succeeds in at least one experiment. Qwen3-32B-finetune achieved a 79.17% sub-task completion rate, outperforming Llama3.1-405B (Paper) (69.05%) by a margin of 10.12%. This margin is significant, particularly when considering that Llama3.1-405B is a much larger model (405B parameters) operating in its native configuration.
@@ 6.2.3 Sub-task Completion Performance (5 Experiments)
In the Access Control (AC) category, Qwen3-32B-finetune achieved a remarkable 14.51%, which is 4.32% higher than Llama3.1-405B (Paper). For Web Security (WS), our model reached 10.91%, outperforming all baselines by a significant margin. Notably, even in categories where task chains are inherently volatile—such as Real-world—the fine-tuned model achieved 21.07%, compared to 17.71% for Llama3.1-405B (Paper) and only 9.78% for Qwen3-32B-base.

### v2

@@ 5.1.2 Victim Environment
    AUTOPENBENCH [9]: Tasks are instantiated as Docker containers, which are hosted on a separate virtual machine to avoid resource contention with the attacker host. This victim VM is configured with 4 vCPUs, 8 GB RAM, and 80 GB storage, and placed in the same NAT network as the attacker machine to ensure direct connectivity.
@@
    AI-Pentest-Benchmark [16]: Vulnerable machines are directly imported from official VulnHub distributions and executed as VMware virtual machines without modification to their default specifications, in order to preserve the original exploitation conditions. All machines are assigned to the same NAT network as the attacker host to guarantee consistent communication.
5.1.3 Evaluation Protocol
Within the unified environment described in Section 5.1.1, xOffense utilizes a core toolchain with specific versions detailed in Table 2. These industry-standard utilities maintain stable command-line interfaces (CLI) and operational syntaxes, ensuring that the agent’s reasoning logic remains functionally compatible with prior benchmark studies despite minor version iterations.
Table 2: Operational toolchain and versions (Kali Linux 2025)
Tool Version
Nmap 7.95
Nikto 2.5.0
WPScan 3.8.28
sqlmap 1.9.2#stable
Metasploit Framework 6.4.50
Hydra 9.5
Enum4linux 0.9.1
Gobuster 3.8.2
Dirb 2.22
The evaluation itself follows a standardized protocol designed to guarantee fair and reproducible comparisons across all evaluated methods. Each agent is executed under identical constraints, including a maximum interaction budget of 5 steps per task, a timeout of 60 minutes per run, and a fixed token budget per model invocation. All methods are restricted to the same set of tools and environment configurations, without access to external resources beyond those explicitly defined.
Each experimental scenario is repeated for 5 independent runs with different random seeds to account for stochasticity in LLM outputs. Reported results are averaged across runs, and we additionally report standard deviations where applicable.
The evaluation follows a goal-oriented setting, where an episode is considered successful if the agent achieves the predefined exploitation objective within the allowed budget. Intermediate actions, such as reconnaissance and vulnerability identification, are not independently scored but contribute to overall task completion.
To prevent potential bias, all models are evaluated on the same subset of benchmark environments, and no task-specific tuning is performed during evaluation. Furthermore, we ensure that training data used for fine-tuning does not overlap with evaluation benchmarks to mitigate data leakage. This protocol is consistently applied to both the proposed framework and all baselines to ensure comparability and reliability of the reported results.
5.2 LLM models
@@
The fine-tuned model, namely Qwen3-32B-finetune is hosted on a dedicated compute node equipped with an NVIDIA A100 GPU having 80 GB VRAM. The same hardware is also used during the fine-tuning process to accelerate training efficiency. For inference, the model is exposed via an API endpoint tunneled through ngrok, allowing the attacker machine to interact with the model as an external service. This separation ensured that LLM inference did not compete with pentest tasks for system resources, while also replicating realistic deployment conditions where models are often served remotely.
5.2.1 Evaluated Baseline Models
The experimental evaluation utilizes a broad spectrum of large language models to establish a rigorous comparative baseline. This selection encompasses leading proprietary models such as GPT-4o, alongside high-performance open-source architectures including Llama3.3-70B, Llama3.1-405B, and DeepSeek-V3. For the mid-scale category, Qwen3-32B-base serves as the foundational open-source representative, which is subsequently optimized into Qwen3-32B-finetune for specialized penetration testing tasks.
Empirical consistency is maintained by enforcing a standardized decoding strategy across all tested models. The inference process is configured with a temperature of 0.5 and a top-p value of 0.9. This specific combination strikes a deliberate balance between the deterministic accuracy required for security tool command synthesis and the generative flexibility necessary for exploring complex attack paths. Furthermore, the top-k parameter is set to 40, allowing for a diverse yet controlled vocabulary selection that prevents the model from generating highly improbable tokens. Each model interaction is allocated a maximum budget of 4096 tokens. This extended context window is essential for processing verbose security logs and generating comprehensive multi-stage exploit scripts without the risk of mid-sentence truncation. This universal hyperparameter setup isolates the architectural design and multi-agent orchestration of each framework as the primary variables driving the observed performance results.
5.2.2 Fine-Tuning Methodology
@@
Unlike conventional large-scale model adaptations, which demand significant computational overhead, xOffense leverages LoRA to achieve domain specialization. LoRA reduces the parameter footprint by freezing the base model’s weights and training only a compact set of adapter matrices. This approach dramatically lowers the number of trainable parameters by over 99%, enabling efficient fine-tuning even on standard GPU infrastructures.
@@
To further address the memory bottlenecks of handling a 32B-parameter model, we employ DeepSpeed ZeRO-3 [30] optimization. ZeRO-3 partitions model states, which includes optimizer, gradients, and parameters across multiple GPUs, achieving linear scalability. Additionally, FlashAttention v2 [4] is integrated to optimize attention computation, reducing memory usage and accelerating training by up to 3x compared to standard attention implementations. These combined techniques allow us to efficiently fine-tune Qwen3-32B for pentest workloads with a significant reduction in hardware demands. And the fine-tuning dataset comprised two main corpora as follows:
@@
    PentestData: This dataset is meticulously curated to encompass domain-specific question-answer pairs, each enriched with synthetically generated CoT reasoning traces. The reasoning steps are encapsulated within <think> tags, enabling the model to learn structured, step-by-step logical deduction processes tailored for pentest scenarios. To construct PentestData, we aggregate and standardize write-ups from over 1,000 machines across leading cybersecurity platforms, including TryHackMe [40], HackTheBox [12], and VulnHub [41]. These write-ups are systematically processed to extract task-specific reasoning paths, exploit procedures, and decision-making sequences relevant to offensive security operations. In addition, we incorporate supplementary pentest datasets from HuggingFace Datasets Hub [14], focusing on cybersecurity knowledge bases, pentest techniques, and practical guides for commonly used security tools. This comprehensive integration ensures that PentestData serves as a robust and diverse resource for training models in autonomous pentest workflows.
@@
    WhiteRabbitNeo: A high-quality JSONL-formatted dataset comprising instruction-response pairs, specifically curated for cybersecurity tasks. Although the original dataset lacked explicit CoT reasoning annotations, it is standardized during preprocessing by appending empty <think> tags to each sample. This structural unification ensures compatibility with CoT-augmented training pipelines and facilitates subsequent fine-tuning for step-by-step reasoning abilities. The dataset draws from real-world offensive and defensive cybersecurity scenarios, encompassing exploitation techniques, payload crafting, and red-team/blue-team interactions, sourced from the WhiteRabbitNeo community contributions [38].
5.2.3 Prompt Setting
To ensure methodological transparency and reproducibility, xOffense adopts a single system-wide initialization prompt (init prompt) that governs the full penetration workflow, rather than using distinct prompt policies for individual agents. This unified prompt defines the global objective, operational boundaries, execution environment assumptions, output contracts, and cross-phase memory usage, thereby enforcing consistent reasoning behavior throughout reconnaissance, scanning, and exploitation.
Concretely, the init prompt specifies four mandatory control dimensions: (i) mission context (target objective and scope constraints), (ii) state continuity (reuse of prior successful actions and shell status), (iii) actionability requirements (tool-compatible and directly executable commands), and (iv) structured output formatting. The unified design reduces prompt-policy drift across phases and improves execution stability when the planner performs iterative update and reflection.
To reduce parsing ambiguity, the init prompt enforces strict output contracts: planning outputs must be serialized within <json></json> tags, while executable commands must be enclosed in <execute></execute> tags. This explicit interface is essential because downstream modules parse model outputs programmatically. In addition, long execution logs are summarized before being re-injected into context, which mitigates context dilution and helps maintain coherent decision-making over multi-step attack trajectories.
The prompt follows a grey-box principle: it does not expose implementation internals, but injects selective operational state, including prior successful or failed tasks, shell continuity indicators, and compressed history summaries. This strategy preserves cross-step coherence without overwhelming the model with raw terminal traces.
For all experiments, prompt execution used a unified inference policy to isolate the effect of architecture and prompting from decoding variance. Specifically, we use temperature =0.5=0.5, top-$p=0.9$, top-$k=40$, and a maximum generation budget of 4096 tokens per interaction, consistent across compared models. In the deployed xOffense implementation, hidden chain-of-thought traces (when produced inside <think> tags) are removed before downstream processing, ensuring evaluation is based only on actionable outputs while retaining stable external behavior.
For reproducibility, a representative structured output schema governed by the init prompt is shown below:
Listing 1: Representative structured planning output used in xOffense prompts.
<json>
"id": "1",
"dependent_task_ids": [],
"instruction": "Enumerate open services on 10.10.10.5 with version detection.",
"action": "Shell"
"id": "2",
"dependent_task_ids": ["1"],
"instruction": "Assess HTTP service on 10.10.10.5:80 for common web vulnerabilities.",
"action": "Web"
</json>
@@ 5.3 Benchmark Datasets
To ensure a rigorous and practically relevant evaluation, we selected two complementary benchmarks that cover both synthetic vulnerabilities and realistic multi-stage exploitation scenarios. Together, these benchmarks provide a balanced testbed for assessing the adaptability and robustness of automated pentest systems.
@@
This benchmark defines a total of 33 pentest tasks, spanning both instructional “in-vitro” exercises and real-world CVEs. The in-vitro set (22 tasks) reflects fundamental vulnerability classes frequently highlighted in industry rankings such as the OWASP Top 10, including weak access control (such as misconfigured sudo, world-writable shadow files), web application flaws (such as path traversal, SQL injection, file upload RCE), and insecure network configurations (such as SNMP misconfiguration, ARP spoofing). In addition, four cryptography tasks evaluate resilience against improper or weak cryptographic implementations, such as brute-forcing Diffie–Hellman keys. Beyond these educational tasks, the benchmark incorporates 11 real-world CVEs ranging from 2014 to 2024 with CVSS scores between 7.5 and 10.0. These include critical vulnerabilities widely recognized for their impact and prevalence, such as Log4Shell (CVE-2021-44228), Heartbleed (CVE-2014-0160), SambaCry (CVE-2017-7494), and Spring4Shell (CVE-2022-22965). By combining foundational categories with critical CVEs, AutoPenBench provides a structured yet realistic environment to evaluate whether agents can transition from basic exploitation to handling high-severity vulnerabilities that have historically dominated real-world attack campaigns.
@@
While AutoPenBench focuses on containerized vulnerabilities, the AI-Pentest-Benchmark evaluates AI agents on complete end-to-end exploitation workflows across 13 real-world vulnerable machines drawn from VulnHub. These machines are categorized by difficulty into easy (such as Victim1, Library2, Funbox, WestWild), medium (such as Cengbox2, Devguru, Symfonos2), and hard (such as Insanity, TempusFugit). Each machine defines a structured set of reconnaissance, exploitation, privilege escalation, and general technique subtasks, amounting to 152 tasks in total. The vulnerabilities embedded within these machines reflect common pentest scenarios, including web application flaws (SQL injection, XSS, CSRF/SSRF), network service weaknesses (FTP/AD enumeration, brute-force authentication), code-level issues (deserialization, command injection), and post-exploitation techniques (cronjob analysis, misconfigured system files, privilege escalation via user access exploitation). Notably, many of these tasks map directly to recurring weakness categories in the CWE Top 25 and OWASP Top 10, ensuring that success on this benchmark corresponds to capabilities relevant in practical offensive security operations. The benchmark is particularly challenging because the ultimate goal is to achieve root access on each machine, requiring coherent reasoning across reconnaissance, exploitation, and privilege escalation stages. Previous studies have shown that even large-scale proprietary models such as GPT-4o and Llama3.1-405B are unable to achieve root-level compromise without human assistance, underscoring the difficulty and realism of this benchmark.
@@ 5.4.1 Overall Task Completion Rate
This metric measures the percentage of target machines successfully compromised like the obtained flag within the allowed interaction budget. Formally:
@@ 5.4.2 Sub-task Completion Rate (1 Experiment)
To gain insight into intermediate stages of pentest, we evaluate sub-task success rates, including service enumeration and vulnerability detection. Each benchmark defines a set of subtasks $S$. A sub-task is considered successful if it is completed in at least one of the five independent runs:
@@ 6.1 Experiment Scenarios
6.2 Task Completion Performance Across pentest Categories
@@ 6.2.1 Overall Task Completion Performance
Table 3 presents the overall task completion rates across all evaluated models on the AutoPenBench dataset. The fine-tuned Qwen3-32B-finetune model achieved a remarkable 72.72% completion rate, substantially outperforming both its base variant, Qwen3-32B-base (30.30%), and other state-of-the-art models, including GPT-4o (21.21%), Llama3.1-405B (Paper) (30.30%), and PentestGPT (9.09%).
The performance disparity between Qwen3-32B-finetune and its base version is particularly noteworthy. Despite having identical model architecture and parameter size (32 billion parameters), the domain-specific fine-tuning enabled a 2.4x improvement in task completion. This validates the effectiveness of our lightweight LoRA fine-tuning pipeline in adapting general-purpose models to specialized pentest workflows.
In the AC category, Qwen3-32B-finetune achieved a 100% success rate, a stark contrast to the 40.00% of Qwen3-32B-base and the 60.00% of Llama3.1-405B (Paper). Similarly, in NS, our model achieved 83.33% completion, surpassing all baselines, including Llama3.3-70B (33.33%) and Qwen3-32B-base (50.00%). Notably, even in the complex Real-world category, Qwen3-32B-finetune attained a 54.54% success rate, outperforming Qwen3-32B-base (27.27%) and PentestGPT (0.00%).
These findings demonstrate that fine-tuning on domain-relevant CoT data and incorporating robust task orchestration mechanisms can enable a quantized, resource-efficient model to match and even surpass the capabilities of larger, general-purpose LLMs in specialized scenarios. The consistent outperformance across categories further validates the robustness of our fine-tuning strategy, especially given the compute-efficient AWQ quantization.
@@ 6.2.2 Sub-task Completion Performance (1 Experiment)
To assess finer-grained capabilities, we evaluated sub-task completion in a single-run experiment (Table 4). The term ”1 Experiment” refers to the overall sub-task completion rate across five experiments, where a sub-task is considered successful if it succeeds in at least one experiment. Qwen3-32B-finetune achieved a 79.17% sub-task completion rate, outperforming Llama3.1-405B (Paper) (69.05%) by a margin of 10.12%. This margin is significant, particularly when considering that Llama3.1-405B is a much larger model (405B parameters) operating in its native configuration.
@@ 6.2.3 Sub-task Completion Performance (5 Experiments)
In the AC category, Qwen3-32B-finetune achieved a remarkable 14.51%, which is 4.32% higher than Llama3.1-405B (Paper). For WS, our model reached 10.91%, outperforming all baselines by a significant margin. Notably, even in categories where task chains are inherently volatile, such as Real-world, the fine-tuned model achieved 21.07%, compared to 17.71% for Llama3.1-405B (Paper) and only 9.78% for Qwen3-32B-base.